this is my question. the measure of the consecutive angles of a parallelogram are 3x-5 & 5x+1.find the measure of each angle of the parallelogram? thank you for helping me to understand this solving problem. tnx
Answer by mananth(16946)


the measure of the angles area 3x-5 & 5x+1.
the opposite angles are equal.
total of all angles= 360 deg.
2(3x-5)+2(5x+1)=360
6x-10+10x+2=360
16x-8=360
16x = 368
x=23 deg.
...
The angles are 3x-5 = 69-5 = 64
5x+1 = 5*23+1=116
The angles are 64 & 116. The opposite two angles being equal.